The algebraic relation d=0.0056s^2+0.14s models he relation between a vehicle's stopping distance d, in meters and it's speed s, in kilometers per hour (km/hr).
What is the stopping distance for a car traveling at 120km/h?

Answer by macston(5194)

.
d=0.0056s^2+0.14s
.
Substitute s=120km/h and solve for d:
.
d=0.0056(120^2)+0.14(120)
d=0.0056(14400)+16.8
d=80.64+16.8=97.44
.
ANSWER: Stopping distance is 97.44 meters.